OpenStudy (amtran_bus):

The graphs in the XY plane of y=x^2 and y=x^2+x+k intersect where x=?

OpenStudy (whpalmer4):

How can you make x^2 = x^2 + x + k?

OpenStudy (amtran_bus):

Substitute?

OpenStudy (whpalmer4):

How about subtracting any terms you see on both sides of the equation...

OpenStudy (whpalmer4):

What does that leave you?

OpenStudy (amtran_bus):

If y=x^2, then just plug that into the longer one to get what you did.

OpenStudy (whpalmer4):

Subtract x^2 from both sides of that equation I gave you, what do you get?

OpenStudy (whpalmer4):

0 = x + k, right? Can you solve that for x?

OpenStudy (amtran_bus):

x+k

OpenStudy (amtran_bus):

X+K=0

OpenStudy (whpalmer4):

Okay, solve for x...

OpenStudy (amtran_bus):

Ohhh.........-K

OpenStudy (whpalmer4):

x =

OpenStudy (whpalmer4):

right! If x = -k, then the two have the same y value.

OpenStudy (amtran_bus):

Thanks and God bless

OpenStudy (whpalmer4):

you'll encounter not a few problems where just making everything match is all you have to do...

OpenStudy (amtran_bus):

Thats right!